OpenStudy (mackenzie2013):

A number divided by 4, minus 3 is 0. What is the number? 0 4 8 12

OpenStudy (calculusxy):

It is 12.

OpenStudy (anonymous):

x/4-3=0..... x/4=3........x=12

OpenStudy (calculusxy):

Since 12/4=3 and 3-3=0.

OpenStudy (anonymous):

Suppose number =x x/4-3=0 x/4=3 x=3x4 x=12

OpenStudy (mackenzie2013):

Thaks !

OpenStudy (mackenzie2013):

Thanks

OpenStudy (calculusxy):

Welcome! Any help, be sure to mention.

OpenStudy (anonymous):